OpenStudy (anonymous):

A 0.430kg soccer ball is initially moving at 2.20m/s . A soccer player kicks the ball, exerting a constant 39.0N force in the same direction as the ball's motion. Over what distance must her foot be in contact with the ball to increase the ball's speed to 5.95m/s ?

OpenStudy (anonymous):

use E=1/2*m*v^2

OpenStudy (anonymous):

use E=1/2*m*v^2 Eball= 1/2*.430*2.20^2= 1.0406 THen find force : M*A=E -> sfa A=E/M 1.0406/.430=2.42 N Since the foot has 39 N and you want 5.95m/s use F*D= 1/2*m*v^2 39*x= 1/2*.430*5.95^2 Solve for x x = .195 I think